OpenStudy (anonymous):

(3-radical 2) (2- radical 8) X and simplify

OpenStudy (anonymous):

Is there an x and an equal sign in there somewhere?

OpenStudy (anonymous):

no theres no equal sign or x. you mulitply them

OpenStudy (anonymous):

its (3-\[\sqrt{2}\] and (2-\[\sqrt{6}\]

OpenStudy (anonymous):

is it sqrt 6 or sqrt 8 ?

myininaya (myininaya):

\[3*2-3*\sqrt{8}-2\sqrt{2}+\sqrt{8}\sqrt{2}=6-3\sqrt{4*2}-2\sqrt{2}+4\] \[=6-6\sqrt{2}-2\sqrt{2}+4=10-8\sqrt{2}\] i hope im right lol

OpenStudy (anonymous):

yes you were thanks myininaya

myininaya (myininaya):

:) yes error free!

OpenStudy (anonymous):

Looks good to me, too